In a meeting a government officer wanted to distribute certain number of apples equally
among the physically handicapped children. If the number of children was 2 less, each of
them would have got one apple more. Again when the ten officers present in the day are
also included, then each of them will get 3 apples less. Find the number of the
handicapped children and the number of apples
1
Expert's answer
2019-12-09T11:29:28-0500

Let a - amount of apples, c - amount of children. Will compose a system of equations describing these two conditions.


"\quad\begin{cases} \cfrac{a}{c-2} - \cfrac{a}{c} = 1 \\
\cfrac{a}{c} - \cfrac{a}{c+10} = 3
\end{cases}"


Will carry out the elementary transformations


"\quad\begin{cases} \cfrac{ac-ac+2a}{c(c-2)} = 1 \\
\cfrac{ac+10a-ac}{c(c+10)} = 3
\end{cases}"


"\quad
\begin{cases} 2a = c(c-2) \qquad\qquad\qquad \small\text{(1)} \\
10a = 3c(c+10) \qquad\qquad\space \small\text{ (2)}
\end{cases}"


Will multiply the first equation by 5, equate the right sides of the equations and determine the "c" .


"\quad 5c(c-2) = 3c(c+10)\\ \quad5c-3c=30+10\\ \quad c=\cfrac{40}2\\ \quad\underline{c=20}"


Will substitute "c" in equation (1) and define the "a"


"\quad2a = 20*(20 - 2)\\ \quad a = \cfrac12*20*18\\ \quad\underline{a = 180}"


Answer: 20 children, 180 apples.